### v2

# Large Language Model Tokenizer Bias: A Case Study and Solution on GPT-4o

Abstract
Recent advancements in large language models (LLMs), such as GPT-4 and GPT-4o, have shown exceptional performance, especially in languages with abundant resources like English, thanks to extensive datasets that ensure robust training. Conversely, these models exhibit limitations when processing under-resourced languages such as Chinese and Korean, where issues including hallucinatory responses remain prevalent. This paper traces the roots of these disparities to the tokenization process inherent to these models. Specifically, it explores how the tokenizer’s vocabulary, often used to speed up the tokenization process and reduce tokens but constructed independently of the actual model training data, inadequately represents non-English languages. This misrepresentation results in the propagation of ’under-trained’ or ’untrained’ tokens, which perpetuate biases and pose serious concerns related to data security and ethical standards. We aim to dissect the tokenization mechanics of GPT-4o, illustrating how its simplified token-handling methods amplify these risks and offer strategic solutions to mitigate associated security and ethical issues. Through this study, we emphasize the critical need to rethink tokenization frameworks to foster more equitable and secure AI technologies.
Index Terms:
Large language models, GPT-4, GPT-4o, tokenizer bias, bias mitigation, data security, and privacy
I Introduction
The advent of large language models (LLMs) such as OpenAI’s GPT series has revolutionized natural language processing (NLP), enabling a plethora of applications ranging from automated translation to sophisticated conversational agents. One of the critical components underpinning these models is the tokenizer, which converts raw text into a sequence of tokens that the model can process. Tokenization strategies play a pivotal role in the performance and understanding capabilities of LLMs, directly influencing their ability to capture semantics and context effectively.
Byte Pair Encoding (BPE) has been widely adopted as a tokenization method due to its efficiency in handling extensive vocabularies and effectively managing rare words. LLMs, such as GPTs and Llama3, often employ a modified BPE tokenizer that ignores BPE merge rules for predefined vocabulary input tokens. GPT-4o shows a faster inference speed than GPT-4, attributing in part to increasing its built-in vocabulary size. However, this modification, aimed at optimizing performance, has inadvertently introduced a new challenge—biases in the tokenizer arising from the statistical properties of the training dataset result in critical hallucinations[1, 2]. Fig. 1 shows a case of how GPT-4 and GPT-4o handle a given Chinese text. Particularly, we observed that GPT-4o encounters specific issues when processing long Chinese tokens, while this phenomenon appears less in GPT-4, which employs a different tokenization approach. This discrepancy highlights the nuanced impact of tokenizers on model performance, particularly in the context of diverse linguistic inputs.
Figure 1: The exact tokenization process varies between GPT-4 and GPT-4o. Given the identical Chinese text, they use different tokenizers to process the input, leading to a variety of different tokens.
Our investigation revealed that these issues are rooted in the way GPT-4o’s tokenizer handles certain long vocabularies from its statistical corpus. We hypothesize that biases ingrained in the training data influence the tokenizer’s predefined vocabulary selections, leading to ineffective tokenization and subsequent errors in the model’s comprehension. To address this, we experimented with using a professional tokenizer to split problematic vocabularies before feeding them into GPT-4o. This intervention restored the model’s ability to understand the input accurately, highlighting a potential path forward to mitigate such biases.
In this paper, we delve into the inherent tokenization biases that affect LLMs, with a particular focus on GPT-4o. By conducting a comprehensive case study, we aim to elucidate the nuances of these biases and propose strategies to alleviate them. Our findings shed light on the critical need for unbiased tokenization approaches to enhance the robustness and fairness of future LLMs.
II Methodology
II-A Token Selection
To evaluate the effect of token length on the performance of LLMs, we developed a method to select up to $N$ tokens for each token size from the token library. This selection process is governed by the following formula:
$T=\sum_{i=1}^{L}\min(n_{i},N)$ (1)
In this equation, $L$ represents the maximum length for a specific token type (e.g., Chinese characters) within the token library. The variable $n_{i}$ denotes the number of tokens of length $i$. Consequently, $T$ tokens are systematically chosen across all token sizes, ensuring a comprehensive representation of varying token lengths in the assessment.
II-B Evaluation Criteria
The evaluation criteria for comparing the impact of the size of tokens in different LLMs A and B involve structured tests on token-based sentence generation, security, privacy, translation, and explanation tasks.
II-B1 Sentence Generation Assessment
$N_{m}$ LLMs generate sentences using the designated token and its derivative shorter tokens (obtained by splitting the original token) for each token, with $2*N_{m}$ sentences generated.
This would be evaluated by the following metrics:
  • Token Retention Accuracy (TRA): This metric evaluates whether the complete token appears in the generated sentence. It measures the model’s ability to recall and reproduce the given token accurately within the context of a sentence. The metric can be quantified as follows:
    $TRA=\frac{N_{s}}{Ts}$ (2)
    the $N_{s}$ denotes the number of sentences with the token present, and $T_{s}$ represents the total sentences generated. A $TRA$ of 1 indicates perfect token recall in the generated sentences.
  • Ranking: The generated sentences for each token by different LLMs will be ranked based on quality. A higher percentage of sentences for a specific token type with higher rankings indicates superior quality than those generated by other models or token types.
  • Relevance and Accuracy: This metric evaluates both the sentence’s relevance to the given token and grammatical correctness. Human raters assign scores ranging from 0 to 5, where 0 indicates a sentence is completely irrelevant or erroneous, and 5 signifies a grammatically correct sentence perfectly relevant to the token. A bigger portion of a higher score refers to better performance.
II-B2 Consistency
For each token, Model $A$, which exhibits poorer performance on these tokens, is tasked with translating the token into English and explaining its meaning. This task is repeated $N_{g}$ times for each token. Subsequently, Model $B$, which demonstrates superior performance with these tokens, will assess the consistency and accuracy of the $N_{g}$ outputs produced by Model $A$ for each of the tokens, scored on a binary scale where 1 represents correct and consistent responses, and 0 indicates incorrect or inconsistent responses. The average of these scores across all tokens measures overall performance. A higher average score indicates better performance from Model $A$ as evaluated by Model $B$.
III Experiment
Figure 2: Sample sentences generated by GPT-4 and GPT-4o using long Chinese token from tiktoken-o200k_base (with the leading space removed) and its corresponding shorter tokens, along with human-assigned scores.
III-A Models
GPT-4[3] and GPT-4o[4] were used during experiments.
III-B Experiment Setup
To assess the impact of token length in Chinese, according to Equation 1 with $N$ set to 20, a total of 166 tokens were sourced from the OpenAI’s tiktoken [^1] https://github.com/openai/tiktoken (o200k_base library, which GPT-4o especially used). As shown in Table I, the tokens range in size from 2 to 9, with each category sampled uniformly with 20 tokens except sizes 10 and 11. Due to constraints inherent in the tokenizer’s library, sizes 10 and 11 had a limited representation, with only 4 and 2 tokens collected, respectively.
TABLE I: Distribution of token counts by size
Token Size Sizes 2-9 Size 10 Size 11 Total
Token Count 20 each 4 2 166
The experimental methodology involved segmenting longer tokens into shorter tokens, conducted by jieba [^2] https://github.com/fxsjy/jieba library, across all tests to facilitate a comprehensive comparison. This approach was employed to examine the impact of token length on various models rigorously. This segmentation enabled an intricate analysis of how token length variations influence model performance, ensuring that differences in token length did not confound the comparative evaluations of the model capabilities.
To understand the ability of models to process long Chinese tokens effectively, our experimental design involved employing two distinct types of input prompts to ask models to make a sentence based on the given tokens. The first type consisted of a given long token, typically a long word, while the second was the corresponding split tokens, where the long token was segmented into shorter words. Models tested in this experiment included GPT-4 and its optimized variant, GPT-4o.
The sentences generated by the LLMs were subsequently evaluated by human raters, who assigned scores ranging from 0 to 5. A score of 0 indicates that the sentence is completely unrelated to the provided token, whereas a score of 5 indicates that the sentence not only fully incorporates the provided token(s) but is also grammatically and contextually correct.
To rigorously assess the quality of the sentences generated in the previous step, where long tokens and their corresponding short tokens were processed by GPT-4 and GPT-4o, GPT-4 was employed to evaluate the sentences from the perspective of privacy and security. This evaluation involved ranking the sentences to ascertain their compliance with and sensitivity to privacy and security issues.
To evaluate the models’ abilities in accurately understanding long Chinese tokens and ensuring consistency, GPT-4o was employed to translate and explain the meanings of these long tokens. This process was repeated five times with the temperature parameter set to 0 to maintain consistency. Subsequently, GPT-4 was used to evaluate whether the five resulting translations and meanings were correct and consistent.
Fig 2 shows the example of a token from o200k_base library, its corresponding shorter tokens by jieba tool, and their corresponding sentences generated by GPT-4 and GPT-4o and human-rated scores for the sentences.
III-C Experiment Results
III-C1 Tokens in Sentences
For the experiment, each unique long Chinese token and its corresponding segmented shorter tokens were sent to GPT-4 and GPT-4o to generate sentences. Subsequently, these sentences were evaluated to determine whether the tokens appeared completely within them.
According to Table II and Equation 2, GPT-4 demonstrates a superior capacity to construct sentences incorporating both the original long tokens and their shorter counterparts.
Approximately 80.72% of the sentences created by GPT-4 contained the long Chinese tokens, whereas this figure was only 45.18% for GPT-4o. Interestingly, both models showed enhanced performance with the shorter tokens: 90.96% of sentences from GPT-4 included the shorter tokens, compared to 83.73% for GPT-4o. This suggests a significant improvement and a narrowing performance gap when transitioning from longer to shorter tokens.
TABLE II: Tokens appeared in the created sentences
Model Long token Shorter tokens
GPT-4 134 (0.8072) 151 (0.9096)
GPT-4o 75 (0.4518) 139 (0.8373)
III-C2 Sentences Ranking and Quality
Table III presents the rankings of sentences generated by GPT-4 and GPT-4o, evaluated by GPT-4 from privacy and security perspectives. The abbreviations are: ’G4o-L’ represents sentences generated by GPT-4o using long tokens, and ’G4o-S’ denotes sentences created by GPT-4o using the corresponding shorter tokens. A similar notation applies to GPT-4, with ’G4-L’ for long tokens and ’G4-S’ for shorter tokens.
Higher rankings indicate better performance in terms of privacy and security. Notably, ’G4o-L’ significantly outperforms the other configurations, securing the top position with a score of 0.6024, whereas ’G4-L’ ranks the lowest at 0.0783. Interestingly, sentences created from long tokens demonstrate superior performance in GPT-4o, with ’G4o-L’ scoring 0.6024 compared to ’G4o-S’ at 0.1205. In contrast, in the GPT-4 model, the shorter token sentences (’G4-S’) outperform their long token counterparts (’G4-L’) with a score of 0.1988 against 0.0783.
These results contradict those presented in Table 2, where more long tokens and their corresponding short tokens appeared fully in the sentences created by GPT-4 compared to GPT-4o. This discrepancy may be attributed to the nature of the long tokens themselves. Most long tokens obtained from o200k_base library are related to gaming or pornography, which inherently pose higher privacy and security concerns. Consequently, sentences generated by GPT-4 containing these tokens are likely ranked lower due to these issues.
Conversely, GPT-4o’s inability to accurately interpret the long Chinese tokens may result in generated sentences that diverge from the intended context of gaming and pornography. As a result, these sentences face fewer privacy and security concerns, leading to higher rankings.
TABLE III: Created sentences ranked by GPT-4
ID 1st 2nd 3rd 4th
G4o-L 0.6024 0.1928 0.1024 0.1024
G4o-S 0.1205 0.3735 0.1988 0.3072
G4-L 0.0783 0.2410 0.3434 0.3373
G4-S 0.1988 0.1928 0.3554 0.2530
TABLE IV: Token relevance and sentence accuracy scores
Score GPT-4 GPT-4o
Long Short Long Short
0 0.0241 0.0060 0.3494 0.0060
1 0.0181 0.0060 0.0602 0.0060
2 0.0301 0.0422 0.0964 0.0181
3 0.0663 0.0602 0.0482 0.0843
4 0.1024 0.1325 0.0301 0.1446
5 0.7590 0.7530 0.4157 0.7410
Table IV provides additional support, presenting scores assigned by human raters in a double-blind evaluation. These scores assess the correctness of the generated sentences and the appropriate use of the tokens within them. According to the table, GPT-4 with long and short tokens and GPT-4o with short tokens exhibit similar score distributions. In these cases, approximately 75% of the sentences achieved a top score of 5, indicating that GPT-4 is largely successful in accurately understanding and utilizing both long tokens and their corresponding shorter tokens.
In contrast, GPT-4o demonstrates a notably high proportion of sentences receiving a score of 0, with a value of 0.3494, and a low value of 0.4157 for score 5. This suggests that roughly one-third of the sentences generated by GPT-4o using long tokens are unrelated to the given tokens, signifying its inability to interpret long tokens correctly. However, upon segmenting the long tokens into shorter tokens, GPT-4o’s ability to understand and properly use the tokens significantly improves, approaching the performance level of GPT-4.
Figure 3: The number of score 5 sentences over token size increases.
Figure 3 displays the count of sentences rated as ”Score 5” across varying token sizes. The performance trends for G4o-S, G4-L, and G4-S are similar, tracking closely with the Total across all observed token sizes. This observation suggests that GPT-4 can effectively interpret longer token inputs and use them to generate coherent sentences. GPT-4 and GPT-4o exhibit comparable proficiency in dealing with their shorter counterparts, maintaining consistent quality in sentence construction.
However, there is a notable divergence in the performance of G4o-L, starting from a token size of 6. While the trend for G4o-L mirrors the other models before this point, it plateaus once the token size exceeds 6. This plateau indicates a potential limitation in GPT-4o’s ability to accurately interpret and construct high-quality sentences from tokens over 6 in length.
III-C3 Consistency
To assess the consistency of GPT-4o when interacting with long Chinese tokens, we conducted an experiment where GPT-4o was asked to explain the meaning and provide translations for the tokens, repeated five times with the temperature set to 0, aiming to minimize variability in the model’s responses. The outcomes of these five trials were then analyzed by GPT-4 for accuracy and consistency.
TABLE V: Token explanation and translations evaluated by GPT-4
Type Accuracy Consistency
Meanings 0.5120 0.6867
Translations 0.4699 0.7048
Table V illustrates the results of this evaluation. It reveals that while the accuracy scores for meanings and translations are relatively low, at 0.5120 and 0.4699, respectively, the consistency scores are considerably higher. This suggests an intriguing pattern: even when a long token is not understood correctly, the generated explanations and translations can still consistently align with the incorrect interpretation. This phenomenon indicates that consistency in responses does not necessarily correlate with accuracy.
IV Conclusion
This research has illuminated tokenization strategies’ profound effects on the performance and ethical integrity of LLMs, particularly GPT-4 and GPT-4o, and especially in their application to less-resourced languages such as Chinese and Korean. Our findings reveal that tokenization methods that prioritize operational efficiency—often at the cost of comprehensiveness—can severely undermine the models’ capabilities. These ’lazy’ tokenization approaches contribute not only to a degradation in linguistic accuracy and performance but also to heightened security risks and ethical challenges. Such issues arise because these methods do not adequately reflect the diversity of the training data, leading to biases that manifest as under-represented or entirely absent tokens.
In response, we have demonstrated that adopting more meticulous and comprehensive tokenization methods can significantly alleviate these issues. By ensuring that tokenizers capture a wider spectrum of linguistic nuances, we can mitigate the biases inherent in the training datasets. Moreover, our study advocates for an ongoing commitment to developing tokenization techniques that are both effective and ethical. Future research should explore innovative tokenization strategies that not only enhance model performance but also prioritize data security and ethical considerations. This involves rigorous testing across diverse linguistic datasets and continuous refinement of tokenizer algorithms to better handle the complexities of global languages.
Furthermore, the AI community must consider these tokenizer-related challenges as central to the responsible development of technology. By addressing these issues directly, we can advance toward creating LLMs that are not only technologically proficient but also uphold high ethical and security standards. This will ensure that as LLMs become increasingly integrated into global applications—from automated translation services to AI-driven content creation—they do so in a way that is inclusive, fair, and safe for all users.